# Subleading correction to the Asian options volatility in the Black-Scholes model

### v1

@@ 1. Introduction
$\Sigma_{\rm LN}^{2}(K,T)$ $=$ $\sigma^{2}\Big{\{}\underbrace{\frac{\log^{2}x}{2J_{\rm BS}(x)}}\underbrace{-\frac{61}{9,\!450}(\sigma^{2}T)+\frac{1}{12}(rT)}$
@@ 3.3. Including a non-zero interest rate $r$
$\Sigma^{2}_{\rm LN}(K,T,S_{0})$ $=$ $\sigma^{2}\Big{\{}\frac{\log^{2}x}{2J_{\rm BS}(x)}-\frac{61}{9,\!450}(\sigma^{2}T)+\frac{1}{12}(rT)$
@@ 3.4. Consistency check and an improved estimate
(78) $v(\rho):=\frac{1}{\rho^{3}}(\rho e^{2\rho}-\frac{3}{2}e^{2\rho}+2e^{\rho}-\frac{1}{2})=\frac{1}{3}+\frac{5}{12}\rho+\frac{17}{60}\rho^{3}+O(\rho^{4})$
@@ 4. Numerical examples
$\hskip 56.9055pt+\,\Big{(}\frac{12,\!073}{1,\!034,\!500}-\frac{5}{525}(\mu+1)\Big{)}x^{2}+O(x^{3})\Big{)}\tau+O(\tau^{2})$
@@
$\Sigma_{\rm LN}^{2}(K,T)$ $=$ $\sigma^{2}\Big{\{}\underbrace{\frac{\log^{2}x}{2J_{\rm BS}(x)}}\underbrace{-\frac{61}{9,\!450}(\sigma^{2}T)+\frac{1}{12}(rT)}$
@@
$\hskip 14.22636pt+\underbrace{\Big{[}\frac{12,\!073}{16,\!632,\!000}(\sigma^{2}T)-\frac{5}{2,\!016}(rT)\Big{]}x^{2}}+O(Tx^{3})+O(T^{2})\Big{\}}\,.$

### v2

@@ 1. Introduction
$\Sigma_{\rm LN}^{2}(K,T)$ $=$ $\sigma^{2}\Big{\{}\underbrace{\frac{x^{2}}{2J_{\rm BS}(e^{x})}}\underbrace{-\frac{61}{9,\!450}(\sigma^{2}T)+\frac{1}{12}(rT)}$
@@ 3.3. Including a non-zero interest rate $r$
$\Sigma^{2}_{\rm LN}(K,T,S_{0})$ $=$ $\sigma^{2}\Big{\{}\frac{x^{2}}{2J_{\rm BS}(e^{x})}-\frac{61}{9,\!450}(\sigma^{2}T)+\frac{1}{12}(rT)$
@@ 3.4. Consistency check and an improved estimate
(78) $v(\rho):=\frac{1}{\rho^{3}}\Big{(}\rho e^{2\rho}-\frac{3}{2}e^{2\rho}+2e^{\rho}-\frac{1}{2}\Big{)}=\frac{1}{3}+\frac{5}{12}\rho+\frac{17}{60}\rho^{3}+O(\rho^{4})$
@@ 4. Numerical examples
$\hskip 56.9055pt+\,\Big{(}\frac{1,\!657}{259,\!875}-\frac{5}{252}(\mu+1)\Big{)}x^{2}+O(x^{3})\Big{)}\tau+O(\tau^{2})$
@@
$\Sigma_{\rm LN}^{2}(K,T)$ $=$ $\sigma^{2}\Big{\{}\underbrace{\frac{x^{2}}{2J_{\rm BS}(e^{x})}}\underbrace{-\frac{61}{9,\!450}(\sigma^{2}T)+\frac{1}{12}(rT)}$
@@
$\hskip 14.22636pt+\underbrace{\Big{[}\frac{1,\!657}{4,\!158,\!000}(\sigma^{2}T)-\frac{5}{2,\!016}(rT)\Big{]}x^{2}}+O(Tx^{3})+O(T^{2})\Big{\}}\,.$